OpenStudy (goformit100):

A force F1 of magnitude 6.00 units acts at the origin in a direction 30.0° above the positive x axis. A second force F2 of magnitude 5.00 units acts at the origin in the direction of the positive y axis. Find graphically the magnitude and direction of the resultant force F1 + F2 .

OpenStudy (goformit100):

@terenzreignz @Eyad

terenzreignz (terenzreignz):

Physics? Free Body Diagram time :D

OpenStudy (goformit100):

Ok ......

terenzreignz (terenzreignz):



terenzreignz (terenzreignz):



terenzreignz (terenzreignz):

So, now, the way to get the resultant vector is to add their respective components :)

OpenStudy (goformit100):

Now my vectors are becoming simpler now

terenzreignz (terenzreignz):

So, you understand from here? If you're given the magnitude and direction (angle) of a vector, you need to break it down into its x and y components... \[\huge \vec v= x\vec i + y\vec j\]

terenzreignz (terenzreignz):

In this case, \[\huge x = |\vec v|\cos (\theta) \\\huge y = |\vec v|\sin(\theta)\]

terenzreignz (terenzreignz):

So, for F1 \[\huge \vec{V1}=[6\cos(\theta)]\vec i + [6\sin(\theta)]\vec j\] NOTE : the i and j here are the horizontal and vertical components of the vectors respectively.

OpenStudy (goformit100):

ok

terenzreignz (terenzreignz):

\[\huge \vec{V1}=[6\cos(30^o)]\vec i + [6\sin(30^o)]\vec j\]

OpenStudy (goformit100):

ok

terenzreignz (terenzreignz):

Sorry, not V1, but F1 \[\huge \vec{F1}=\left[3\sqrt3\right]\vec i + \left[3\right]\vec j\]

terenzreignz (terenzreignz):

Now, if you do F2, the magnitude is 5, and the direction (angle) is 90 degrees... \[\huge \vec{F2}=[5\cos(90^o)]\vec i +[5\sin(90^o)]\vec j\]\[\huge \vec{F2}=[0]\vec i + [5]\vec j\]

terenzreignz (terenzreignz):

So... \[\Large \vec{F1}+\vec{F2}=(3\sqrt3) \ \vec{i}+3\vec j + 0\vec i+ 5\vec j\]\[\huge \vec{F1}+\vec{F2}= \left(3\sqrt 3\right)\vec i +8\vec j\]

terenzreignz (terenzreignz):

\[\huge \vec{F1}+\vec{F2}=\left<3\sqrt3 \ , \ 8\right>\] Now, find magnitude and direction :)

OpenStudy (goformit100):

Thank now i can solve it :)

terenzreignz (terenzreignz):

Yes, and now, (hopefully) you can now solve similar problems :D